OpenStudy (anonymous):

The measure of two of the interior angles of a triangle are 68° and 70°. What is the measure of the third interior angle?

OpenStudy (anonymous):

68 + 70 = 148 180 - 148 = 32 interior angles of triangle total 180 degrees

OpenStudy (anonymous):

the answers are 222 42 22 and 62

OpenStudy (anonymous):

was just a mistake in the calculation there. the sum of angles of a triangle is 180° so, if you add up the other two angles (68 + 70), you have 138 then subtract that from the total of 180: 180 - 138 = 42°

OpenStudy (anonymous):

thanks altig :)